# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 7.2 Relation to Abdulkadiroğlu and Sönmez (1999)
Upon discovering this highly intriguing potential application in healthcare, Ünver and I expressed to Roth our keen interest in initiating a joint research project on kidney exchange. To our disappointment, Roth initially declined our invitation, citing two reasons. First, he was skeptical that economists could independently write the paper without the involvement of a transplantation expert. Second, he believed there wasn’t much formal analysis left to undertake in kidney exchange, as this setting was a direct application of Abdulkadiroğlu and Sönmez (1999).
@@
At first glance, Roth’s reservations seemed justified. The parallel with Abdulkadiroğlu and Sönmez (1999), forming the basis of the ”discovery-invention chain” that links innovations in house allocation to kidney exchange, is grounded in the following observations.
@@
As explored in Section 5.6.1, Abdulkadiroğlu and Sönmez (1999) introduced a generalized top trading cycles (TTC) mechanism, allocating two types of houses to two types of individuals. Existing tenants, one type of individuals, have the option to retain their current houses if they choose to do so. The other type of individuals, the newcomers, do not have any claim to specific houses. The two types of houses include the occupied houses earmarked for their existing tenants and the vacant houses which are not tied to any specific individuals.
How does this setting relate to an application in healthcare?
From a modeling perspective, in the context of kidney exchange:
@@
  • Patients with incompatible (or less than ideal) living donors play a role similar to existing tenants who desire to receive more-preferred houses than the houses they currently occupy.
  • Living donors of patients play a role similar to occupied houses by existing tenants.
  • Patients in DD lists who have no living donors play a role similar to newcomers who do not currently occupy any house.
@@
  • DD kidneys or living-donor kidneys from non-directed donors play a role similar to vacant houses. [^72] In Roth et al. (2004), the analogy with vacant houses was only given for DD kidneys. The same analogy was later introduced for non-directed donor kidneys in Roth et al. (2006), which also introduced and advocated for non-simultaneous donor chains initiated by non-directed (a.k.a. good Samaritan) donors.
Similar to existing tenants who can retain their current houses in Abdulkadiroğlu and Sönmez (1999), patients with living donors can retain their donors if they wish to do so. Just as newcomers have no claim on any specific house, patients on the DD list also have no claim on any specific kidney. Finally, as in the case where all individuals collectively own all vacant houses in house allocation, all patients have joint claims to all DD kidneys.
Drawing on these analogies, the Abdulkadiroğlu and Sönmez (1999) variant of the TTC mechanism finds a direct application in kidney exchange. By managing claims for “unattached” houses, which could be either vacant or vacated during the procedure, through an exogenous priority list, the TTC mechanism organizes two distinct types of transactions: cycles and chains.
In a cycle, existing tenants exchange their occupied houses, with paired kidney exchange corresponding to a cycle involving only two patients with living donors. In a chain, an individual exchanges their priority for an “unattached” house, and the remaining individuals trade their currently occupied houses. List exchange corresponds to a chain involving one patient with a living donor and the highest-priority patient on the DD list.
Despite the strong parallels between kidney exchange and house allocation with existing tenant settings, Ünver and I were convinced of the substantial value in pursuing this project. One key reason for our optimism was the ethical concerns surrounding list exchanges and the potential to modify the Abdulkadiroğlu and Sönmez (1999) variant of the TTC mechanism to address these issues. While using exogenous priority lists, as in TTC, remained a viable approach for regulating chains in kidney exchange, we recognized that alternative chain selection rules could mitigate, or even eliminate, the adverse effects of list exchange on blood-type O patients on the DD list.
We believed this added flexibility would be highly valuable in this context. For example, welfare gains from chains, which represent generalized list exchanges, could be incorporated into the system without negatively impacting the balance of blood type O patients on the DD list. By leveraging this insight, both goals of the transplantation community could be addressed through a generalization of the TTC mechanism, rather than treating kidney exchange simply as a direct application of Abdulkadiroğlu and Sönmez (1999). In our view, this generalization offered significant practical and theoretical value.
Another reason for our optimism was the potential to simulate the number of transplants that could be achieved through an organized system. After studying numerous papers on kidney transplantation, we became confident that this project could be undertaken by economists alone. We shared our thoughts with Roth and suggested that he might join us later, after reviewing the first draft.
Roth gave us his blessing to proceed if we were determined. Ünver and I spent additional time reading on the topic and began addressing the formal and computational aspects of the model. In terms of the fit between the real-world application and its formulation in theoretical research, kidney exchange posed a far more challenging case than school choice. Not only did kidney exchange involve far more details and considerations from more distant fields (e.g., transplantation, immunology, and medical ethics), but there was often no consensus on many aspects of the problem in those respective literatures. [^73] For example, as thoroughly discussed in Roth et al. (2004), the literature was unclear on whether a reduction in HLA mismatch between the patient and donor would increase graft survival or not. Therefore, from the beginning, it was apparent that there was no single best model for kidney exchange.
Despite these challenges, we finalized the first draft of our paper within a few months.
@@ 7.3 First Kidney Exchange Model in Roth, Sönmez, and Ünver (2004)
As a use-inspired theory project, our approach to kidney exchange differed markedly from my earlier approach to school choice. In school choice, we offered a set of refined policy recommendations (see Section 6.10). In contrast, at this stage, our focus in kidney exchange was primarily on providing insights into the virtues of:
  1. 1.
    Establishing a database for patient-donor pairs, and
  2. 2.
    Organizing kidney exchanges through a systematic mechanism that enhances both utilitarianism and equity—objectives highlighted in the transplantation literature.
Our efforts paid off, leading Roth to join our project upon reviewing the first draft. Shortly thereafter, in September 2003, we circulated our analysis in Roth et al. (2003), and a year later, we published a more streamlined version in Roth et al. (2004).
Positioned as the next innovation in a “discovery-invention cycle”—building on earlier inventions by medical professionals, in particular paired kidney exchange and list exchange, and progressing through the foundation laid by Abdulkadiroğlu and Sönmez (1999)—our paper conveyed a clear message: The number of transplants facilitated by kidney exchanges can be increased by several orders of magnitude while simultaneously avoiding harm to the pool of blood-type O patients on the DD list.
@@
In essence, we demonstrated that significant improvements can be achieved in both primary objectives emphasized in the preceding kidney exchange literature in medicine: utilitarianism and equity. These improvements were obtained by developing advanced versions of earlier kidney exchange inventions from this literature and optimizing their implementation. Within this minimalist exercise, the primary root causes of the shortcomings in existing kidney exchange programs were twofold: (1) the absence of a comprehensive database containing living donor information, and (2) the lack of a systematic mechanism to organize these exchanges. A secondary root cause was the limitation of both types of kidney exchanges to their most basic forms, involving the smallest possible number of patient-donor pairs—two for paired kidney exchange and one for list exchange. We demonstrated that the effectiveness of a kidney exchange program could be significantly increased by allowing for 3-way or larger direct exchanges or adopting larger donor chains initiated by volunteers of list exchanges.
By addressing these concerns with a range of conceptual and technical innovations in Roth et al. (2003, 2004), we provided members of the transplantation community with both a methodology and a class of mechanisms that offer significant improvements in both utilitarianism and equity—the two principles emphasized in the Consensus Statement in Abecassis et al. (2000). As we would soon find out, the timing of our paper was perfect.
@@ 7.4 Initial Phase of Kidney Exchange in New England
[…]
Since it involved only a single patient-donor pair, arranging a list exchange might have been expected to be much easier. However, the eligibility criteria for this more controversial form of kidney exchange included ensuring that no paired kidney exchange—the preferred option between the two—was feasible with any other patient registered across all 14 transplant centers in New England.
@@ 7.5 Forging a Partnership with New England Medical Professionals
As far as modeling itself is concerned, formulating a new kidney exchange model to address Delmonico’s reservations was relatively straightforward. However, the analysis of the resulting model presented a completely different story. At that time, neither Ünver nor I were familiar with the formal techniques needed to analyze this new model.
@@
This challenge could have posed a serious problem if basic theory had not come to our rescue. Fortunately, just a few months earlier, economists Anna Bogomolnaia and Hervè Moulin had published a paper on a closely related theoretical model, albeit in a more restrictive setting with an assumption that precluded its direct adoption for kidney exchange (Bogomolnaia and Moulin, 2004). If only we could analyze the more general version of the model we needed, we could then address Delmonico’s reservations. Building on the foundational work of Gallai (1963, 1964) and Edmonds (1965, 1971) in discrete optimization, this is exactly what we did in Roth et al. (2005b).
@@
The progress we made with our second kidney exchange model was well-received by Delmonico and his colleagues in the New England transplantation community. This breakthrough laid the groundwork for establishing the New England Program for Kidney Exchange (NEPKE), founded by our team of three design economists in collaboration with doctors Francis Delmonico and Susan Saidman (Roth et al., 2005a). Approved by the Renal Transplant Oversight Committee of New England in September 2004, NEPKE became the world’s first organized kidney exchange system to utilize formal techniques from market design and optimization.
@@
Starting with the Alliance for Paired Donation in 2006, many kidney exchange clearinghouses worldwide followed NEPKE’s example in the subsequent years, transforming the way living-donor kidney transplants are organized in numerous countries. [^76] See, for example, Manlove and O’Malley (2012); Anderson et al. (2015); Veale et al. (2017); Furian et al. (2019, 2020); Biro et al. (2019, 2021); Salman et al. (2023); Duus Weinreich et al. (2023); Yilmaz et al. (2023, 2024) for living-donor organ exchange programs and techniques where market design experts took active roles. In that sense, the establishment of NEPKE stands out as one of the most consequential policy achievements of my career.
@@
This interaction with Delmonico left me with lasting impressions for two additional reasons. From a methodological perspective, it provided firsthand insight into the importance of custom-made theory in persuading decision makers and experts in other fields of the value of our contributions. As elaborated in this monograph, my approach of seeking policy impact through use-inspired theory, closely aligned with practical realities, has become my primary methodology for numerous applications of market design in the years that followed. In my view, this approach inherently incorporates a natural and effective persuasion strategy within minimalist market design.
@@
The final reason for this lasting impression lies in how this interaction influenced the global evolution of kidney exchange, shaping my future research and policy efforts in living-donor organ exchange. Currently, kidney exchange programs worldwide predominantly restrict participation to incompatible patient-donor pairs, resulting in substantial global welfare losses. This restriction poses a particular challenge, especially for blood type O patients with donors of blood types A, B, or AB. The issue arises because, aside from tissue-type incompatibility, blood type O donors are medically compatible with their co-registered patients.
@@
In a notable departure from this approach, Roth et al. (2005b)—developed at the request of Delmonico and fostering our collaboration—assumes that patients are indifferent among all compatible donors. Unfortunately, this modeling choice resulted in the loss of the intrinsic mechanism present in Roth et al. (2003, 2004), which incentivized compatible pairs to participate in kidney exchange. Since NEPKE served as a blueprint for kidney exchange programs subsequently launched in the U.S. and many other countries, this led to a substantial global welfare loss relative to our innovation’s true potential.
To incentivize blood-type compatible pairs to participate in kidney exchange programs, various policies have been proposed in the academic literature or implemented in the field since then. Before examining two of these policies in Sections 7.7 and 7.8, let’s explore how our partnership with New England medical professionals paved the way for a few other breakthroughs in kidney exchange.

### v4

@@ 7.2 Relation to Abdulkadiroğlu and Sönmez (1999)
In the early 2000s, Ünver and I were both faculty members at Koç University in İstanbul. As discussed in Section 5.6.2, my first project with him focused on the house allocation with existing tenants model (see Sections 5.45.6). During the 2002–2003 academic year, while Ünver was visiting Roth at Harvard University, we realized that this model had a direct application in healthcare—kidney exchange. Discovering this connection between our work and a healthcare application was exhilarating; never before had economic theory research shown such clear potential to save lives. Motivated by this promise, the three of us soon began a joint research project on kidney exchange, aiming both to advance the emerging field of market design and to influence policy.
Yet pursuing this line of work was far from straightforward, and our initial enthusiasm was tempered by two immediate concerns. The first was whether economists could undertake such a project independently, without the involvement of a transplantation expert. The second, and more pressing, was whether meaningful formal analysis remained to be done in kidney exchange, given its close connection to Abdulkadiroğlu and Sönmez (1999) and the HA–TTC (or, equivalently, the YRMH–IGYT) mechanism introduced there as a solution.
After reviewing numerous papers on kidney transplantation, we became confident that economists could pursue this project on their own. The second concern, however, proved more substantive, especially from a pure research perspective.
@@
The parallel with Abdulkadiroğlu and Sönmez (1999), forming the basis of the “discovery–invention chain” that links innovations in house allocation to kidney exchange, rests on the following observations.
@@
As discussed in Section 5.6.1, the HA–TTC mechanism allocates two types of houses to two types of individuals. Existing tenants, one type of individual, have the option to retain their current houses if they so choose. The other type, the newcomers, have no claim to any specific house. The two types of houses are the occupied houses earmarked for their existing tenants and the vacant houses, which are not tied to any specific individual.
This structure has a natural counterpart in kidney exchange, where patients and donor kidneys play roles analogous to tenants and houses:
@@
  • A patient with an incompatible (or less-than-ideal) living donor is analogous to an existing tenant seeking a more-preferred house than the one they currently occupy.
  • The kidney from a patient’s living donor is analogous to an occupied house (i.e., the tenant’s endowment).
  • A patient on the DD list with no living donor is analogous to a newcomer who does not currently occupy any house.
@@
  • A DD kidney or a kidney from a non-directed living donor is analogous to a vacant house. [^82] In Roth et al. (2004), this analogy with vacant houses was drawn only for DD kidneys. The same analogy was later extended to kidneys from non-directed donors in Roth et al. (2006), which also introduced and advocated non-simultaneous donor chains initiated by non-directed (a.k.a. Good Samaritan) donors.
Similar to an existing tenant who may retain their current house, a patient with a living donor may keep their own donor’s kidney if they so choose. Just as a newcomer has no claim to any particular house, a patient on the DD list has no claim to any particular kidney. Finally, just as a vacant house is not tied to any individual, a DD kidney is not tied to any specific patient.
Drawing on these analogies, the HA–TTC mechanism finds a direct application in kidney exchange. Figure 19 illustrates the isomorphism between the two settings. By managing claims to “unattached” house—whether vacant or vacated during the procedure—through an exogenous priority list, the HA–TTC mechanism organizes two distinct types of transactions: cycles and chains. Figure 18 provides an example of both. [^83] Strictly speaking, by employing a technical construct that treats individuals and houses as separate entities, the HA–TTC algorithm transforms each chain of individuals in YRMH–IGYT into a cycle of individuals and houses. To distinguish between the two types of cycles in HA–TTC, we adopt the terminology from its YRMH–IGYT equivalent, using chain nomenclature for the latter.
Figure 18: Kidney exchange cycle (left) and kidney exchange chain (right), each involving three patients with living donors. In the cycle, the donor of P1 donates kidney K1 to P2, the donor of P2 donates K2 to P3, and the donor of P3 donates K3 to P1. In the chain, the donor of P1 donates K1 to P2 in exchange for P1 gaining priority on the deceased-donor list. This initiates a sequence: the donor of P2 donates K2 to P3, the donor of P3 donates K3 to L (a patient on the deceased-donor list), and with this upgraded priority P1 ultimately receives kidney D from the deceased-donor pool.
In a cycle, individuals with endowments (e.g., existing tenants or patients with living donors) exchange them among themselves. In a chain, an individual gives up priority for an unattached object to receive another’s endowment. That person then either ends the chain by taking an unattached object or passes their endowment to another individual, allowing the chain to continue. The process repeats until the final individual, whose endowment was taken by the previous participant, receives an unattached object. Paired kidney exchange corresponds to a cycle involving two patients with living donors, while list exchange corresponds to a chain involving one patient with a living donor and the highest-priority patient on the DD list.
Figure 19: Isomorphism between the house allocation with existing tenants model and kidney exchange. House allocation with existing tenants: T1 and T2 represent existing tenants endowed with occupied houses H1 and H2, respectively. N represents a newcomer with no endowment, and V represents a vacant house unattached to any individual. Kidney exchange: P1 and P2 represent patients endowed with kidneys K1 and K2, respectively, from their living donors. L represents a patient who has no living donor, is on the deceased-donor list only, and has the highest priority for kidney K1 when it is assigned through the deceased-donor list. D represents the deceased-donor (or non-directed donor) kidney that is assigned to patient P1 in exchange for kidney K1 from P1’s living donor, which is donated to the deceased-donor list. Abbreviations: LD: Living donor; $D$: Deceased donor; NDD: Non-directed donor; PKE: Paired Kidney Exchange.
@@ 7.3 First Kidney Exchange Model in Roth, Sönmez, and Ünver (2004)
Though kidney exchange closely mirrors house allocation with existing tenants, we saw clear value in pursuing this project. Our optimism stemmed both from the prospect of demonstrating substantial welfare gains through computational simulations of how many additional transplants an organized system could enable and from the ethical concerns surrounding list exchanges and the possibility of modifying the HA–TTC mechanism to address them. Guided by the principles emphasized in the Consensus Statement, we began exploring how such a modification could better align with the ethical priorities of the transplantation community.
Top Trading Cycles and Chains (TTCC) Mechanism.
While the use of an exogenous priority list, as in HA–TTC, remained a viable way to regulate chains in kidney exchange, we found that alternative chain selection rules could better serve the community. We therefore developed a new mechanism, TTCC. [^84] See Table 3 for a comparison of TTCC with its antecedents—G–TTC, HA–TTC, and SC–TTC—in terms of their property-rights structures, the types of trades they generate, and the tools they employ.
Table 3: Comparison of Top Trading Cycles (TTC) variants: G–TTC in housing markets, HA–TTC in house allocation with existing tenants, SC–TTC in school choice, and TTCC in kidney exchange. Organized by property-rights structures and the induced trade types.
G–TTC HA–TTC SC–TTC TTCC
Model Housing markets House allocation w/ existing tenants School choice Kidney exchange
Object types House OH & AH School seat LD & DD
Property rights Pvt. ownership
Pvt. ownership (OH)
Priority list (AH)
Priority list
Pvt. ownership (LD)
Priority list ($D$)
Trade types House $\rightarrow$ house
OH $\rightarrow$ OH
OH $\rightarrow$ AH-priority
Priority $\rightarrow$ priority
LD $\rightarrow$ LD
LD $\rightarrow$ $D$-priority
  • Abbreviations: Pvt.: Private; OH: Occupied House; AH: Available House; LD: Living-Donor; DD: Deceased-Donor.
Building on HA–TTC, the TTCC mechanism allows for larger cycles and chains than those in paired kidney exchange or list exchange, offering important efficiency gains (see Figure 18 for a depiction of a cycle and a chain, each involving three patients with living donors). Its flexibility in chain selection provides a valuable policy lever: it allows welfare gains from chains to be incorporated without undermining the balance of blood-type O patients on the DD list. In this way, it could mitigate—or even eliminate—the adverse effects of list exchanges on these patients.
Thus, rather than treating kidney exchange as a direct application of the HA–TTC mechanism from Abdulkadiroğlu and Sönmez (1999), TTCC leverages this flexibility to advance both efficiency and equity. The more we learned from the kidney exchange and living-donor liver transplantation literatures, the more confident we became that TTCC offers significant practical and theoretical value.
@@
This theoretical advance also suggested a practical path forward. At the time, existing kidney exchange programs in the early 2000s suffered from three main root causes: (1) the absence of a comprehensive database containing living-donor information, (2) the lack of a systematic mechanism to organize exchanges, and (3) the restriction of both paired and list exchanges to their most basic forms, involving only the smallest possible number of patient–donor pairs (two for paired exchange and one for list exchange). Through TTCC, conceived as a minimalist intervention, we showed that addressing these root causes could yield significant improvements in utilitarianism and equity—the two primary objectives of the transplantation community, highlighted in the Consensus Statement.
We circulated our analysis in Roth et al. (2003) in September 2003 and, a year later, published a more streamlined version in Roth et al. (2004). [^85] In terms of the fit between the real-world problem and its formulation in theoretical research, kidney exchange posed a far more challenging application than school choice. It involved many more details and considerations from diverse fields (e.g., transplantation, immunology, and medical ethics), and there was often no consensus on key aspects of the problem. For example, as thoroughly discussed in Roth et al. (2004), the literature was unclear on whether a reduction in HLA mismatch between patient and donor would improve graft survival. From the outset, therefore, it was clear that there could be no single best model for kidney exchange. Positioned as the next innovation in a “discovery—invention cycle”—building on earlier medical innovations such as paired kidney exchange and list exchange, and on the foundation laid by Abdulkadiroğlu and Sönmez (1999)—our work provided the transplantation community with both a methodology and a mechanism, TTCC, that embodied the principles emphasized in the Consensus Statement.
As we would soon learn, the timing of our paper was perfect.
@@ 7.4 Initial Phase of Kidney Exchange in New England
In its early stages, the program did not establish a unified database of willing living donors for kidney patients that was accessible to all transplant centers involved. Consequently, from 2001 to 2004, the program arranged only 5 paired kidney exchanges, providing kidney transplants to a total of 10 patients (Roth et al., 2005b).
Since it involved only a single patient–donor pair, arranging a list exchange might have been expected to be much easier. However, the eligibility criteria for this more controversial form of kidney exchange included ensuring that no paired kidney exchange—the preferred option between the two—was feasible with any other patient registered across all 14 transplant centers in New England.
@@ 7.5 Forging a Partnership with New England Medical Professionals
As far as modeling itself is concerned, formulating a new kidney exchange model to address Delmonico’s reservations was relatively straightforward. However, the analysis of the resulting model presented a completely different story. At that time, none of us were familiar with the formal techniques required to analyze this new model.
@@
This challenge could have posed a serious problem had basic theory not come to our rescue. Fortunately, just a few months earlier, economists Anna Bogomolnaia and Hervé Moulin had published a paper on a closely related theoretical model, albeit in a more restrictive setting that precluded its direct adoption for kidney exchange (Bogomolnaia and Moulin, 2004). [^87] In Bogomolnaia and Moulin (2004), agents are partitioned into two groups, and each agent can be matched only with an agent on the other side of the economy. In our setting, by contrast, any patient can be matched with any other patient to exchange their donors. If only we could analyze the more general version of the model we needed, we could then address Delmonico’s reservations. Building on the foundational work of Gallai (1963, 1964) and Edmonds (1965, 1971) in discrete optimization, this is exactly what we did in Roth et al. (2005b).
@@
The progress we made with our second kidney exchange model was well-received by Delmonico and his colleagues in the New England transplantation community. This breakthrough laid the groundwork for establishing the New England Program for Kidney Exchange (NEPKE), founded by our team of three economic designers in collaboration with doctors Francis Delmonico and Susan Saidman (Roth et al., 2005a). Approved by the Renal Transplant Oversight Committee of New England in September 2004, NEPKE became the world’s first organized kidney exchange system to utilize formal techniques from market design and optimization.
@@
Starting with the Alliance for Paired Donation in 2006, many kidney exchange clearinghouses worldwide followed NEPKE’s example in the subsequent years, transforming the way living-donor kidney transplants are organized in numerous countries. [^88] See, for example, Manlove and O’Malley (2012); Anderson et al. (2015); Bingaman et al. (2017); Veale et al. (2017); Furian et al. (2019, 2020); Biro et al. (2019, 2021); Salman et al. (2023); Duus Weinreich et al. (2023); Yilmaz et al. (2023, 2024) for living-donor organ exchange programs and techniques where market design experts took active roles. In that sense, the establishment of NEPKE stands out as one of the most consequential policy achievements of my career.
@@
This interaction with Delmonico left me with lasting impressions for two additional reasons. From a methodological perspective, it provided firsthand insight into the importance of custom-made theory in persuading decision-makers and experts in other fields of the value of our contributions. As elaborated in this monograph, my pursuit of policy impact through use-inspired theory, closely aligned with practical realities, has become my primary methodology for numerous applications of market design in the years that followed. In my view, this methodology naturally embeds an effective persuasion strategy within minimalist market design, going a long way toward satisfying the framework’s third pillar.
@@
The final reason for this lasting impression lies in how this interaction influenced the global evolution of kidney exchange, shaping my future research and policy efforts in living-donor organ exchange. Currently, kidney exchange programs worldwide predominantly restrict participation to incompatible patient–donor pairs, resulting in substantial global welfare losses. This restriction poses a particular challenge, especially for blood type O patients with donors of blood types A, B, or AB. The issue arises because, aside from tissue-type incompatibility, blood type O donors are medically compatible with their patients.
@@
In a notable departure from this approach, Roth et al. (2005b)—developed at the request of Delmonico and marking the beginning of our collaboration—assumes that patients are indifferent among all compatible donors. Unfortunately, this modeling choice eliminated the intrinsic mechanism present in Roth et al. (2003, 2004) that incentivized compatible pairs to participate in kidney exchange. Since NEPKE served as a blueprint for kidney exchange programs subsequently launched in the U.S. and many other countries, this led to a substantial global welfare loss relative to the innovation’s true potential. In the case of the U.S., Sönmez et al. (2020) estimates that including blood-type compatible pairs in kidney exchange could increase the number of transplants by as much as 160%.
To incentivize blood-type compatible pairs to participate in kidney exchange programs, various policies have been proposed in the academic literature or implemented in the field since then. Before examining one of these policies in Section 7.7, let’s explore how our partnership with New England medical professionals paved the way for a few other breakthroughs in kidney exchange.